# Gibbs randomness-compression proposition: An efficient deep learning

### v2

@@ Abstract
A proposition that connects randomness and compression is put forward via Gibbs entropy over set of measurement vectors associated with a compression process. The proposition states that a lossy compression process is equivalent to directed randomness that preserves information content. The proposition originated from the observed behaviour in newly proposed Dual Tomographic Compression (DTC) compress-train framework. This is akin to tomographic reconstruction of layer weight matrices via building compressed sensed projections, via so-called weight rays. This tomographic approach is applied to previous and next layers in a dual fashion, that triggers neuronal-level pruning. This novel model compress-train scheme appears in iterative fashion and acts as a smart neural architecture search, The experiments demonstrated the utility of this dual-tomography producing state-of-the-art performance with efficient compression during training, accelerating and supporting lottery ticket hypothesis. However, random compress-train iterations having similar performance demonstrated the connection between randomness and compression from statistical physics perspective, we formulated the so-called Gibbs randomness-compression proposition, signifying randomness-compression relationship via Gibbs entropy. Practically, the DTC framework provides a promising approach for massively energy- and resource-efficient deep learning training.
Index Terms:
@@ I Introduction
Connections between randomness and data compression are noted earlier within the algorithmic information framework of Solomonoff-Kolmogorov-Chaitin [1, 2, 3] (SKC). In summary these works aim at reaching a definition of randomness from model selection perspective [4], i.e., algorithmic entropy or a complexity measure. Striking connections to algorithmic entropy and Gibbs entropy are also explored from fundamental conceptual point of view [5, 6, 7]. Statistical mechanics connections and information are established between Gibbs and Boltzmann entropies [8, 9, 10, 11]. However, relationship among compression process, entropy, information and randomness is always made quite generic in SKC framework. This manifestation was by design quite generic without strict mathematical condition, i.e., aimed at finding the shortest algorithm or explanation that can generate the random sequence. This is similar in entropy production studies [12].
In this work, we propose a mathematical condition among these concepts in a more formal manner inspired by newly proposed deep learning compress-train framework. We describe a technique called inverse compressed-sensing in dual fashion, Dual Tomographic Compression (DTC) compress-train for deep learning in an iterative manner as an example of Gibbs randomness-compression proposition. Such interplay between physics and deep machine learning is now well-established [13, 14, 15, 16, 17, 18].
II Background
@@
Our purpose is establishing a mathematical relationship between randomness and complexity, whereby complexity appears as a compression process. In this direction, we leverage a couple of distinct research areas, randomness and complexity, compressed sensing and neural network architecture compression.
@@
The definition of randomness [19] and quality of generating random numbers [20] appears to be connected subjects within the complexity, compressing data [6], empirically [21] and SKC algorithmic randomness context. This framework commonly known as Kolmogorov Complexity (KC), algorithmic complexity or algorithmic randomness and its applications are vast [22, 23]. KC or SKC framework, makes a generic statement about information content of sequence $s$ of binary digits of length $N$, $K(s)$, defined to be the size of the smallest program in bits that generates $s$, the sequence. The connection to compression is implicit here and its mathematical definition to compression process appears as only a single compression step. It is hinted that there is a positive correlation between compression ratio and randomness [22]. Moreover, the interest of randomness in this manner also explored from biological evolution stand point [24], which is more relevant in our case due to proposed train-compress framework having mutation-like process.
@@
Compressed sensing (CS) framework has shown huge leap in statistical signal processing and information theory as it shows one can reconstruct an unknown signal with much less measurements under sparsity constraints [25, 26, 27, 28, 29]. CS involves random sampling and it provides an excellent tool to investigate the relationship between compression and randomness.
@@
Reducing size of neural networks has attracted research interest since their inception, such as earlier works on Hopfiled networks [30] in statistical physics community. This pioneering work identifies clipping as removing existing connections and pruning re-learning, i.e., fine-tuning after clipping. Similar clipping approach first applied in computer science by LeCun et al. [31], fun title of brain damage is used, reflecting the performance degradation due to clipping, hinting at magnitude pruning, so called small saliency that small weights don’t affect the performance that much, also recently for semi-supervised learning [32, 33]. With the advent of deep learning [34, 35] and generative models producing ever increasing size large deep learning architectures due to attention [36] and diffusion [37] mechanisms, necessitates producing much smaller deep learning models for both training and inference efficiency, as energy consumption and ethical use is raised as the most important aspect in industry and academic practice [38] and sparse model size reduction post-model training became important [39], for dynamic cases as well [40, 41]. Lottery ticket hypothesis [42], provides a deeper perspective in compressing or finding smaller deep learning architectures, via magnitude pruning, i.e., brain damage, stating that subnetworks are possible that give similar performance as the initial deep architecture due to initialization chance, i.e., a lottery ticket. Applying the lottery ticket approach, iteratively pruning during training, leads to train-compress procedure, a compression process, which is the main example tool for our quest for finding relationship between randomness and compression. Such process of train-compress as a meta algorithm has been already proposed [43, 44], before lottery tickets were proposed. Effectiveness of random only pruning also reported recently [45]. Using compressed sensing in pruning recurrent networks is also explored, however it was only on to reduce size of independently appearing network components not on connections [46]. Using deep learning within compressed sensing framework is also explored recently in a different angle [47, 48].
The contributions of the paper are presented in the following sections Inverse Compressed Sensing, Dual Tomographic Compression and the main proposition, Gibbs randomness-compression proposition. These are supported by in-depth state-of-the-art (SoTA) numerical experiments on canonical vision deep learning test system.
@@
III Inverse Compressed Sensing
We apply compressed sensing on the learned weights. However, as weights are known, instead of measuring the signal (weights), we do have it. This implies so called inverse compressed sensing, not to be confused with inverse problem nature of the original compressed sensing framework. Inverse here implies, we construct an hypothetical measurement process from known signal, and try to reconstruct the signals sparse version via conventional compressed sensing, i.e.,so called weight rays. In case of deep learning layers, total weights associated to neuron’s input and output will be used instead of single CS, dual nature of tomographic process. Essentially, we found a sparse projection of the learned weights, which might be sparser due to compressed sensing framework. In the notation, we will use weights, instead of calling a signal vector [28]. Feedforward fully-connected weights are assumed, however generalizing to other layer types should be obvious.
Given, learned weights vector $w\in\mathbb{R}^{m}$, that is flatten from fully connected layer matrix, representing all neurons state at a given learning time. Our purpose is to find sparse projection of $w$, $w_{s}\in\mathbb{R}^{n}$ where by $n$ is defined via sparsity level $s$, $n=m\cdot s$. Within CS framework, we define two matrices, one measurement matrix $\Psi\in\mathbb{R}^{mxm}$, here discrete cosine transform matrix and randomization matrix, here normal Gaussian, $\Phi\in\mathbb{R}^{nxm}$. Then CS matrix reads $\Theta=\Phi\cdot\Psi\in\mathbb{R}^{nxm}$. The inverse bit as we discusses, comes to play when we generate a hypothetical measurement vector, $y=\Phi\cdot w$ as a measurement vector. Then we minimize the following unconstrained version of CS procedure, to reconstruct the learned weights, so called weight ray $w_{r}$,
@@ III Inverse Compressed Sensing
IV Dual Tomographic Compression: Neuronal-level pruning
@@
V Gibbs randomness-compression proposition
@@
The proposition is inspired by DTC’s formulation in the previous section. Given a compression process of $n$ steps on a vector $w$, it progresses with measurement vectors $y_{i}$, $i=1,...,n$ in an iterative fashion with a random mechanism ${\bf R}$, as each step corresponds to a compression ratio $s_{i}$ and a fitness value, i.e., performance metric $f_{i}$. The sizes of measurement vectors are monotonically decreasing as the compression ratio compares to the initial signal. A normalized distribution of measurements is formed and its Gibbs entropy is computed as $G_{i}$, i.e., from normalized histogram bins of $k$, where $G_{i}=-\Sigma_{k}p_{k}log_{2}p_{k}$.
@@
The Gibbs randomness-compression proposition states that there is an almost perfect relationship between the two functions $(s_{i},f_{i})$ and $(s_{i},G_{i})$. This proposition establishes that directed randomness, because of its iterative nature, within the compression process is related to compression performance. As compression with preserved fitness is related to the level of complexity, the proposition establishes a relationship between randomness and complexity. In other words, a lossy compression process is equivalent to directed randomness that preserves information content, as evidenced by the retention of performance or fitness.
@@
VI Numerical results
We have embedded DTC into a canonical vision task, classifying MNIST dataset [49] via softmax output. Basic deep learning models are implemented with hidden layer size of 512 neurons as starting point for train-compress iterative pruning. A fixed seed Xavier initialization [50] is used for reproducibility. We apply dual compressed sensing at every 200 batches of size 512, triggering inverse compressed sensing procedure using SCS algorithm [51] for minimizing for weight rays, for simplicity we didn’t use bias terms. We utilize pytorch [52] for deep learning and cvxpy [53] for DTC steps.
TABLE I: Test accuracy for different pruning approaches over different sparsity levels in iterative pruning.
Sparsity Magnitude Prune Random Prune DTC Prune
@@ VI Numerical results
We record the test accuracy of the reduced architecture after each train-compress cycle for different pruning approaches: magnitude only, random and DTC based magnitute pruning. As shown in Table 1, Figure 1 and Figure 2. DTC and random pruning perform superior. We observe that DTC shows superior performance against random pruning after 85 percent sparsity. For very high sparsity levels after 95 percent, randomness shows a bit better performance but this might be due to very small network that has quite fluctuating results. Depending on requirements very good performance is also retained even after 95 percent of the neurons are pruned.

### v3

@@ Abstract
A proposition that connects randomness and compression is put forward via Gibbs entropy over set of measurement vectors associated with a compression process. The proposition states that a lossy compression process is equivalent to directed randomness that preserves information content. The proposition originated from the observed behavior in newly proposed Dual Tomographic Compression (DTC) compress-train framework. This is akin to tomographic reconstruction of layer weight matrices via building compressed sensed projections, via so-called weight rays. This tomographic approach is applied to previous and next layers in a dual fashion, that triggers neuronal-level pruning. This novel model compress-train scheme appears in iterative fashion and acts as a smart neural architecture search: also called compression aware training. The experiments demonstrated the utility of this dual-tomography during training: method accelerates and supports lottery ticket hypothesis. However, random compress-train iterations having similar performance demonstrated the connection between randomness and compression from statistical physics perspective, we formulated the so-called Gibbs randomness-compression proposition, signifying randomness-compression relationship via Gibbs entropy. The proposition is supported with the experimental evidence, resulting in very high correlation between learning performance vs. Gibbs entropy over compression ratios. Practically, the DTC framework provides a promising approach for massively energy- and resource-efficient deep learning training.
keywords:
@@ 1 Introduction
Connections between randomness and data compression are noted earlier within the algorithmic information framework of Solomonoff-Kolmogorov-Chaitin [solomonoff64, kolmogorov65, chaitin66] (SKC). In summary these works aim at reaching a definition of randomness from model selection perspective [chaitin75], i.e., algorithmic entropy or a complexity measure. Striking connections to algorithmic entropy and Gibbs entropy are also explored from fundamental conceptual point of view [bennett82, zurek89, machta99a]. Statistical mechanics connections and information are established between Gibbs and Boltzmann entropies [jaynes65a, jaynes57a, buonsante16, rondoni00]. However, relationship among compression process, entropy, information and randomness is always made quite generic in SKC framework. This manifestation was by design quite generic without strict mathematical condition, i.e., aimed at finding the shortest algorithm or explanation that can generate the random sequence. This is similar in entropy production studies [suzen25a]. In this work, we propose a mathematical condition among these concepts in a more formal manner inspired by newly proposed deep learning compress-train framework. We describe a technique called inverse compressed-sensing in dual fashion, Dual Tomographic Compression (DTC) compress-train for deep learning in an iterative manner as an example of Gibbs randomness-compression proposition. Such interplay between physics and deep machine learning is now well-established [sherrington, hopfield82, ackley85, suzen22c, suzen2023h, suzen24].
2 Background
@@
In order to establish a mathematical relationship between randomness and complexity, where the concept of complexity appears as a compression process, a concrete cases are indispensable. In this direction, we leverage a couple of distinct research areas, randomness and complexity, compressed sensing and neural network architecture compression, as a foundational background of the work.
@@
The definition of randomness [compagner91] and quality of generating random numbers [oneill14] appears to be connected subjects within the complexity, compressing data [zurek89], empirically [chang10] and SKC algorithmic randomness context. This framework commonly known as Kolmogorov Complexity (KC), algorithmic complexity or algorithmic randomness and its applications are vast [li08, zurek91]. KC or SKC framework, makes a generic statement about information content of sequence $s$ of binary digits of length $N$, $K(s)$, defined to be the size of the smallest program in bits that generates $s$, the sequence. The connection to compression is implicit here and its mathematical definition to compression process appears as only a single compression step. It is hinted that there is a positive correlation between compression ratio and randomness [li08]. Moreover, the interest of randomness in this manner also explored from biological evolution stand point [wagner12], which is more relevant in our case due to proposed train-compress framework having mutation-like process.
@@
Compressed sensing (CS) framework has shown huge leap in statistical signal processing and information theory as it shows one can reconstruct an unknown signal with much less measurements under sparsity constraints [donoho06a, candes06, candes06tao, baraniuk07, eldar12a]. CS involves random sampling and it provides an excellent tool to investigate the relationship between compression and randomness.
@@
Reducing size of neural networks has attracted research interest since their inception, such as earlier works on Hopfield networks [janowsky89] in statistical physics community. This pioneering work identifies clipping as removing existing connections and pruning re-learning, i.e., fine-tuning after clipping. Similar clipping approach first applied in computer science by LeCun et al. [lecun89], fun title of brain damage is used, reflecting the performance degradation due to clipping, hinting at magnitude pruning, so called small saliency that small weights don’t affect the performance that much, also recently for semi-supervised learning [shwartz2024compress, shani2025tokens]. With the advent of deep learning [lecun15a, schmidhuber15a] and generative models producing ever increasing size large deep learning architectures due to attention [vaswani17] and diffusion [sohl15] mechanisms, necessitates producing much smaller deep learning models for both training and inference efficiency, as energy consumption and ethical use is raised as the most important aspect in industry and academic practice [menghani23, zhu25] and sparse model size reduction post-model training became important [hoefler2021sparsity], for dynamic cases as well [han2021dynamic, tatarnikova2025optimization]. Lottery ticket hypothesis [frankle18], provides a deeper perspective in compressing or finding smaller deep learning architectures, via magnitude pruning, i.e., brain damage, stating that subnetworks are possible that give similar performance as the initial deep architecture due to initialization chance, i.e., a lottery ticket. Applying the lottery ticket approach, iteratively pruning during training, leads to train-compress procedure, a compression process, which is the main example tool for our quest for finding relationship between randomness and compression. Such process of train-compress (or compression aware training) as a meta algorithm has been already proposed [han15, carreira18, zimmer25], before lottery tickets were proposed. Effectiveness of random only pruning also reported recently [gadhikar23]. Using compressed sensing in pruning recurrent networks is also explored, however it was only on to reduce size of independently appearing network components not on connections [zhen21a]. Using deep learning within compressed sensing framework is also explored recently in a different angle [wu2019deep, machidon2023deep]. The contributions of the paper are presented in the following sections Inverse Compressed Sensing, Dual Tomographic Compression and the main proposition, Gibbs randomness-compression proposition. These are supported by in-depth learning task on canonical vision deep learning test system as a baseline experimental evidence.
@@
3 Inverse Compressed Sensing
We apply compressed sensing on the learned weights. However, as weights are known, instead of measuring the signal (weights), we do have it. This implies so called inverse compressed sensing, not to be confused with inverse problem nature of the original compressed sensing framework. Inverse here implies, we construct an hypothetical measurement process from known signal, and try to reconstruct the signals sparse version via conventional compressed sensing, i.e.,so called weight rays. In case of deep learning layers, total weights associated to neuron’s input and output will be used instead of single CS, dual nature of tomographic process. Essentially, we found a sparse projection of the learned weights, which might be sparser due to compressed sensing framework. In the notation, we will use weights, instead of calling a signal vector [baraniuk07].
Given, learned weights vector $w\in\mathbb{R}^{m}$, that is flatten from any layer matrix without any dependency on the layer type, representing all neurons state at a given learning time. Our purpose is to find sparse projection of $w$, $w_{s}\in\mathbb{R}^{n}$ where by $n$ is defined via sparsity level $s$, $n=m\cdot s$. Within CS framework, we define two matrices, one measurement matrix $\Psi\in\mathbb{R}^{mxm}$, here discrete cosine transform matrix and randomization matrix, here normal Gaussian, $\Phi\in\mathbb{R}^{nxm}$. Then CS matrix reads $\Theta=\Phi\cdot\Psi\in\mathbb{R}^{nxm}$. The inverse bit as we discusses, comes to play when we generate a hypothetical measurement vector, $y=\Phi\cdot w$ as a measurement vector. Then we minimize the following unconstrained version of CS procedure, to reconstruct the learned weights, so called weight ray $w_{r}$,
@@ 3 Inverse Compressed Sensing
4 Dual Tomographic Compression: Neuronal-level pruning
@@
4.1 Interactions between consecutive reconstructions
Here, we take one of the simplest interactions as $w_{r}^{p}+w_{r}^{q}$ for clipping neurons in a ranked manner. However, as this is a framework more complex interactions could be developed further but our purpose here is to establish the DTC’s utility as a framework.
4.2 Layer types and propagation over layers
We express the framework for fully-connected layers but DTC doesn’t depends on the layer types within the neural network. Only requirement is mapping layer into weight rays, i.e., mapping incoming and outgoing connections into vectors. This can be achieved trivially by flattening any tensor representations of the layers to a one dimensional vector for DTC procedure, such as convolution or transformer architectures.
DTC should be applied to each hidden-layer back-to back, so called in forward propagating the DTC. Not to be confused with backpropagation. This is a simple propagation of the DTC procedure. In the case of multiple hidden-layers, DTC procedure becomes easier as incoming connections would already pruned due to clipped neurons from the previous layer.
We present this train-compress cycles in Algorithm 1 as an overall summary. Algorithm description aimed at presenting a representation of the framework with its logical components, rather than implementation details, in order to convey the core idea.
Input: NN number of learn-compress cycles, $L$ hidden-layers with $j$ as index, $n_{j}$ number of neurons at layer $j$, weight rays $w_{j}^{p}$ and $w_{j}^{q}$, $s_{i}$ target rate of prunning (decreasing monotonically)
Output: $L$ hidden layers with reduced size
Data: Training set $X_{tr}$, testing set $X_{te}$
/* Learn-compress cycles: Within batch learning, smaller network at each batches */
1 for i=1,,Ni=1,...,N do
2   Train the network on the given batch(s) taken from $X_{tr}$
 /* Layer-wise compression: DTC propagating over layers. */
3 for j=1,,Lj=1,...,L do
4      Inverse Compressed sensing:
5      Compute measurement matrices $\Theta$ and measurements vector $y$.
6      Dual $\ell_{1}$ minimization, find weights rays $w_{j}^{p}$ and $w_{j}^{q}$.
7      Clipping the layer (neurons) based on quantile at $s_{i}$ percent.
8      Ranking is achieved with using $w_{j}^{p}+w_{j}^{q}$ in order.
9      Compress the layer: Update layers with kept neurons, updating $n_{j}$
10
 /* For L >1 take mean of all Gibbs entropies */
11   Compute Gibbs entropy $G_{i}(s_{i})$ over measurement vectors $y_{j}$.
12   Compute the performance $f_{i}(s_{i})$ of the compressed network on the test set $X_{te}$
Algorithm 1 An algorithmic summary of Dual Tomographic Compression (DTC) based train-compress cycles. This is a representation of the general approach in the framework.
4.3 Computational Complexity
An assessment of the computational complexity of applying DTC follows. The primary compute overhead due to DTC are optimization cycles. This burden is compensated by the train-compress approach. Because in this framework we don’t need to complete entire training before applying the compression, recall compression aware training. We quantify this by the number of optimization calls over cycles against if we were to run full training and compression. $M$ cycles of compression and $N$ batches for learning and compressions, number of Layers $L$.
  • 1.
    DTC with train-compress
    Order of $N\cdot L$ calls to optimizers both learning and compression in batch learning. Note that in this case $N=M$ as we embed the compression within the training loop.
  • 2.
    Full train compress without DTC
    Order of $N\cdot M$ calls to optimizers for learning and assuming compression wouldn’t need an optimizer call.
We have deduced that, even though DTC brings some optimization load, computational complexity is compensated in a manner that, DTC framework is less compute heavy compare to full training and compress frameworks comparatively as $M$ compression cycles are usually much larger than the number of layers $L$ due to gradual compression, i.e., because compressing in large reductions in small steps are quite unlikely.
5 Gibbs randomness-compression proposition
@@
The proposition is inspired by DTC’s formulation in the previous section. Given a compression process of $n$ steps on a vector $w$, it progresses with measurement vectors $y_{i}$, i=1,,ni=1,...,n in an iterative fashion with a random mechanism $\bf R$}, as each step corresponds to a compression ratio $s_{i}$ and a fitness value, i.e., performance metric $f_{i}$. The sizes of measurement vectors are monotonically decreasing as the compression ratio compares to the initial signal. A normalized distribution of measurements is formed and its Gibbs entropy is computed as $G_{i}$, i.e., from normalized histogram bins of $k$, where $G_{i}=-\Sigma_{k}p_{k}log_{2}p_{k}$.
@@
The Gibbs randomness-compression proposition states that there is an almost perfect relationship between the two functions $(s_{i},f_{i})$ and $(s_{i},G_{i})$, denoted by $f_{i}(s_{i})$ and $G_{i}(s_{i})$. This proposition establishes that directed randomness, because of its iterative nature, within the compression process is related to compression performance. As compression with preserved fitness is related to the level of complexity, the proposition establishes a relationship between randomness and complexity. In other words, a lossy compression process is equivalent to directed randomness that preserves information content, as evidenced by the retention of performance or fitness.
5.1 Analysis of the proposition
More formally, we express the idea in terms of relationship between two comonotonic functions, being highly correlated.
Theorem 5.1 (Gibbs-randomness proposition).
Given sequential lossy compression process consists of $M$ consecutive compression cycle. At each compression cycle reduces the data size $s_{i}$ percent, representing model size, with relative to uncompressed size, $i=1,..,M$, where $s_{0}=1.0$. At each compression cycle, a measure vector ${\bf y}_{i}$ due to a randomized algorithm with associated Gibbs entropy $G_{i}(s_{i})$ and a corresponding learning performance $f_{i}(s_{i})$ after the compression cycle i.e., performance on the reduced model, are generated. If learning performance degradation and entropy reduction due to compression kept small in consecutive cycles $|f_{i}-f_{i-1}|<\epsilon_{i}$ and $|G_{i}-G_{i-1}|<\delta_{i}$ during entire learn-compress cycles, then functions $f_{i}$ and $G_{i}$ are highly correlated.
Proof.
A logical proof follows, functions $f_{i}(s_{i})$ and $G_{i}(s_{i})$ are non-increasing monotonically (comonotonic) due to quasi-static decreases $\epsilon_{i}$ and $\delta_{i}$ locally, while mean differences $\Delta$ are close to zero, $\Delta(\epsilon_{i})=\Delta(\delta_{i})\rightarrow 0$ as $|s_{i}-s_{i-1}|\rightarrow 0$. These functions are comoving over compress-learn cycle (functions evolve synchronously), hence are highly correlated. ∎
We known that functions $f_{i}(s_{i})$ and $G_{i}(s_{i})$ are mechanistically connected due to fact that measurement vector ${\bf y}_{i}$ is defined on the network weights, and learned weights directly effect the network learning performance. Because of random process, here Dual Tomographic Compression (DTC) generates ${\bf y}_{i}$, hence the theorem establishes the connection between randomness and compression.
The theorem is expressed for a single hidden-layer but it is readily applicable to entire neural network by simplify repeating it to all hidden-layers back to back at a given compress-train cycle. It can be applied to any layer type by mapping the weight structure to measurement vector ${\bf y}_{i}$. This is demonstrated in the Algorithm 1 in detail.
@@ 5.1 Analysis of the proposition
6 Experiments and discussions
We have embedded DTC into a canonical vision task, classifying MNIST dataset [lecun02] via softmax output. A deep learning model is implemented with hidden layer size of 512 neurons as starting point for train-compress iterative pruning.
We apply dual compressed sensing at every 200 batches of size 512, triggering inverse compressed sensing procedure using SCS algorithm [odono16] for minimizing for weight rays, for simplicity we didn’t use bias terms. The selection of 200 batches were made such that performance is increasing sufficiently, even it is gradual, and limited by the trade-off between compute budget and compressing in a meaningful rate.
Hyperparameter and related details include Adam optimizer’s learning rate of $0.001$, ReLUReLU activations, fixed Xavier initialization [kumar17a] , batch normalization for the baseline network architecture with fully connected layers. We utilize PyTorch [paszke17] for deep learning and cvxpy [diamond16] for DTC steps.
In the conducted experiments, two more different clipping approach is implement for fair comparison. Namely random and magnitude pruning. Random clipping is just randomly drops neurons at the specified sparsity levels. In the case of magnitude pruning we simply rank the highest weight inputs to given neurons. In total we run three compress-train cycles to compare the test performances, as summarized in Table 1.
Table 1: Test accuracy for different pruning approaches over different sparsity levels in iterative pruning.
Sparsity Magnitude Random DTC
@@ 6 Experiments and discussions
6.1 Learning performance over compression cycles
We record the test accuracy of the reduced architecture after each train-compress cycle for different pruning approaches: magnitude only, random and DTC based magnitude pruning. As shown in Table 1, Figure 1 and Figure 2. DTC and random pruning perform superior. We observe that DTC shows a good performance even against random pruning after 85 percent sparsity. However, random pruning is shown remarkable resilience. But we have clearly shown that DTC outperforms magnitude pruning.
For very high sparsity levels after 95 percent, randomness shows a bit better performance. This isn’t related to network initialization, a fixed Xavier was used making results on the table fully repeatable. While compression process is robust to initialization that performance fluctuations are negligible, rather this is due to fact that after 95 percent compression network remain with handful of neurons in the hidden layer. Depending on requirements very good performance is also retained even after 95 percent of the neurons are pruned.